How will the velocity of a gas molecule vary if its molecular weight is 1. Under which of the following volumes will 1.00 mol of an ideal gas exhibit the greatest pressure at 300 K? A. 0.01 L B. 0.10 L C. 1.00 L D. 10.0 L
Using ideal gas law to find volume
1
Expert's answer
2019-09-06T04:17:35-0400

"v= \sqrt{\dfrac{3×8.314×300}{0.001×6.02×10^{23}}}=1.242×10^{-17} m/sec-molecule"

Boyle's law: "P\alpha 1/V"

Max pressure- V=0.01L

Answer: A. 0.01L